OpenStudy (anonymous):

15/30 in lowest terms = 6/12 in lowest terms =

OpenStudy (anonymous):

1/2 1/2

OpenStudy (calculusxy):

all divisible by the numerator

OpenStudy (mathstudent55):

Find the largest number that you can divide the numerator and denominator by. Then divide them by that number.

OpenStudy (calculusxy):

15/15=1 30/15=2 = 1/2 6/6=1 12/6=2 = 1/2